# Interpreting Graph Neural Networks with In-Distributed Proxies

### v1

@@ 6.2 Quantitative Evaluation (RQ1)
Figure 4: Visualization of explanation results from different explanation models on BA-2motifs. The generated explanations are highlighted by bold black edges.
@@
Figure 5: Visualization of explanation results from different explanation models on Benzene. The generated explanations are highlighted by bold orange edges.
The results demonstrate that ProxyExplainer provides the most accurate explanations across all datasets. Specifically, it improves the AUC scores by an average of $7.5\%$ on synthetic datasets and $12.3\%$ on real-world datasets. Comparisons with baseline methods highlight the advantages of our proposed explanation framework. Besides, ProxyExplainer captures underlying explanatory factors consistently across diverse datasets. For instance, MixupExplainer exhibits proficiency on the synthetic BA-2motifs dataset but performs poorly on the real-world Benzene dataset. The reason is that MixupExplainer relies on the independence assumption of explanation and non-explanation subgraphs, which may not hold in real-world datasets. In contrast, ProxyExplainer consistently demonstrates high performance across different datasets, showcasing its robustness and adaptability. We also replace the GCN with the Graph Isomorphism Network (GIN) in Appendix D.3 to demonstrate that ProxyExplainer can explain different GNN models.
Metric BA-2motifs BA-3motifs Alkane-Carbonyl
GT PGE Proxy GT PGE Proxy GT PGE Proxy
Deg. 0.846 0.785 0.032 0.646 0.209 0.135 0.663 0.525 0.225
Clus. 0.601 0.684 0.319 0.834 0.741 0.625 0.051 0.051 0.051
Spec. 0.278 0.289 0.023 0.259 0.108 0.103 0.635 0.469 0.049
Sum. 1.725 1.759 0.374 1.740 1.059 0.864 1.349 1.046 0.325
Metric Benzene Fluoride-Carbonyl MUTAG
GT PGE Proxy GT PGE Proxy GT PGE Proxy
Deg. 0.412 0.500 0.279 0.546 0.478 0.275 0.598 0.526 0.235
Clus. 0.059 0.001 0.052 0.049 0.049 0.049 0.029 0.029 0.071
Spec. 0.278 0.198 0.096 0.385 0.349 0.075 0.442 0.421 0.141
Sum. 0.749 0.698 0.427 0.980 0.876 0.399 1.070 0.976 0.448
6.3 Alleviating Distribution Shifts (RQ2)
In this section, we assess ProxyExplainer’s ability to generate in-distribution proxy graphs. Due to the intractable of direct computation, we follow the previous work (Chen et al., 2023) to compare distributions of multiple graph statistics, including degree distributions, clustering coefficients, and spectrum distributions, between the generated proxy graphs and original graphs. Specifically, we utilize Gaussian Earth Mover’s Distance kernel when computing MMDs. Smaller MMD values indicate similar graph distributions. For comparison, we also include the Ground truth explanations and the ones generated by PGExplainer.
The results are shown in Table 2. “GT” denotes the MMDs between the ground truth explanations and original graphs. “PGE” represents the MMDs between explanations generated by PGExplainer and original graphs. “Proxy” denotes the MMDs between proxy graphs in our method and original graphs. We have the following observations, first, the MMDs between Ground truth and original graphs are usually large, verifying our motivation that a model trained on original graphs may not have correct predictions on the OOD explanation subgraphs. Second, the explanations generated by a representative work, PGExplainer, are often OOD from original graphs, indicating that the original GIB-based objective function may be sub-optimal. Third, in most cases, proxy graphs generated by our method are with smaller MMDs, demonstrating their in-distribution property.
6.4 Ablation Studies (RQ3)
In this section, we conduct ablation studies to investigate the roles of different components. Specifically, we consider the following variants of ProxyExplainer: (1) w/o $G^{\Delta}$: in this variant, we remove the non-explanatory subgraph generator (VGAE), which is the bottom half as shown in Figure 2; (2) w/o $\mathcal{L}_{\text{KL}}$: in this variant, we remove the KL divergence from the training loss in ProxyExplainer; (3) w/o $\mathcal{L}_{\text{dist}}$: in this variant, we remove the distribution loss from ProxyExplainer. The results of the ablation study on BA-2motifs and Alkane-Carbonyl are reported in Figure 3.
Figure 3 illustrates a notable performance drop for all variants, indicating that each component contributes positively to the effectiveness of ProxyExplainer. Especially, in the real-life dataset, Alkane-Carbonyl, without the in-distribution constraint, w/o $\mathcal{L}_{\text{dist}}$ is much worse than ProxyExplainer, indicating the vital role of in-distributed proxy graphs in our framework. Extensive ablation studies on other datasets can be found in the Appendix D.2.
@@ 6.5 Case Studies
In this part, we conduct case studies to qualitatively compare the performances of ProxyExplainer against others. We adopt examples from BA-2motifs and Benzene in this part. We show visualization results in Figure 4 and Figure 5. Explanations are highlighted with bold black and bold orange edges, respectively. From the results, our ProxyExplainer stands out by generating more compelling explanations compared to baselines. Specifically, ProxyExplainer maintains clarity without introducing irrelevant edges and exhibits more concise results compared to alternative methods. The visualized performance underscores ProxyExplainer’s ability to provide meaningful and focused subgraph explanations. More visualizations on these two datasets can be found in the Appendix D.5.
@@ Appendix A Notations
Symbols
Descriptions
$\mathcal{G}$
A set of graphs
$G,{\mathcal{V}},{\mathcal{E}}$
Graph instance, node set, edge set
$v_{i}$
The $i$-th node
${\bm{X}}$
Node feature matrix
${\bm{A}}$
Adjacency matrix
${\bm{Z}}$
Node representation matrix
$Y$
Label of graph $G$
$\mathcal{C}$
A set of labels
$G^{*}$
Optimal explanatory subgraph
${\mathcal{G}}^{*}$
A set of $G^{*}$
$G^{\prime}$
Candidate explanatory subgraph
$G^{\Delta}$
Non-explanatory graph
$\tilde{G}$
Proxy graph of $G^{\prime}$ with a fixed distribution
h, $\textit{{h}}^{\prime}$, $\tilde{\textit{{h}}}$
Graph embedding
$d$
Dimension of node feature
$D$
Dimension of node latent embedding
$f(\cdot)$
To-be-explained GNN model
$\Psi_{\psi}$
Explanation function
$\psi$
Parameter of the explanation function
$t_{\psi}(\cdot)$
Perturbation function
$P_{\mathcal{G}}$
Distribution of original training graphs
$P_{{\mathcal{G}}^{\prime}}$
Distribution of explanation subgraphs
$Q_{\bm{\phi}}$
Parameterized function of $P(\tilde{G}|G^{\prime})$
$\bm{\phi}$
Model parameters of $Q_{\bm{\phi}}$
$\bm{\phi}^{*}$
Optimal $\bm{\phi}$
$\alpha$
Balance parameter between $I(G,G^{\prime})$ and $I(Y,G^{\prime})$
$\bar{{\mathcal{E}}}$
The set of node pairs that are unconnected in $G$
$\tilde{p}_{uv}$
Probability of node pair $(u,v)$ in $\tilde{G}$
$\beta$
A hyper-parameter to get a trade-off between connected and unconnected node pairs
$f_{\text{enc}}(\cdot)$
The front part of GNNs that learns node representations
$f_{\text{cls}}(\cdot)$
The back part of GNNs that predicts graph labels based on node embeddings
$\sigma(\cdot)$
Sigmoid function
$\tau$
Temperature hyper-parameter for approximation
$\lambda$
A hyper-parameter in Proxy loss function
$\mathcal{L}_{\text{dist}}$
Distribution loss between $\tilde{G}$ and $G$
$\mathcal{L}_{\text{KL}}$
KL divergence between distribution of ${\bm{Z}}^{\Delta}$ and its prior
$\mathcal{L}_{\text{proxy}}$
Proxy loss
$\mathcal{L}_{\text{exp}}$
Explainer loss
@@ Appendix B Algorithm
           Update parameters in the explainer with backpropagation.
@@ C.1 Datasets
BA-3motifs (Chen et al., 2023). BA-3motifs is an extended dataset inspired by the BA-2motifs and contains 3,000 synthetic graphs. Each base graph is accompanied by one of three different patterns: house, cycle, or grid.
Alkane-Carbonyl (Sánchez-Lengeling et al., 2020). The Alkane-Carbonyl dataset consists of a total of 4,326 molecule graphs that are categorized into two different classes. Positive samples refer to molecules that have both alkane and carbonyl functional groups. The ground-truth explanation includes alkane and carbonyl functional groups in a given molecule.
Benzene (Sánchez-Lengeling et al., 2020). Benzene consists 12,000 molecular graphs from the ZINC15(Sterling & Irwin, 2015) database, which can be classified into two classes. The main goal is to determine if a Benzene ring is exited in each molecule. In cases where multiple Benzene rings are present, each Benzene ring is treated as a distinct explanation.
Fluoride-Carbonyl (Sánchez-Lengeling et al., 2020). The Fluoride-Carbonyl dataset has 8,671 molecular graphs. The ground-truth explanation is based on the particular combination of fluoride atoms and carbonyl functional groups present in each molecule.
Table 4: Statistics of molecule datasets for graph classifcation with ground-truth explanations.
BA-2motifs
BA-3motifs
Alkane-Carbonyl
Benzene
Fluoride-Carbonyl
MUTAG
Graphs
1,000
3,000
4,326
12,000
8,671
4,337
Average nodes
25.00
21.92
21.13
20.58
21.36
29.15
Average edges
25.48
29.51
44.95
43.65
45.37
60.83
Original graph
Features
None
None
Atom
Atom
Atom
Atom
Ground truth explanation
House, cycle
House,cycle, grid
Alkane,C=O
Benzene Ring
${}^{-}$,C=O
[…]
@@
BA-2motifs
BA-3motifs
Alkane-Carbonyl
Benzene
Fluoride-Carbonyl
MUTAG
Train Acc
0.999
0.997
0.979
0.930
0.951
0.850
Val Acc
1.0
0.997
0.986
0.923
0.956
0.834
Test Acc
1.0
0.977
0.975
0.915
0.951
0.804
@@ C.2 Baselines
To evaluate our model, we well-train the GCN model in to ensure that it takes good performance in graph classification tasks. The results are displayed in Table 5. We utilize a well-established explanation method as a benchmark baseline PGExplainer. Furthermore, for a comprehensive comparison, we incorporate various post-hoc explanation methods, including GradCAM, GNNExplainer, PGExplainer, ReFine, and MixupExplainer.
  • [leftmargin=*]
@@ D.1 Extensive Distribution Analysis
As shown in previous work (Zhang et al., 2023b), another way to approximate the distribution differences of graphs is to compare their vector embeddings. Here, we adopt the same setting and use Cosine similarity and Euclidean distance to intuitively approximate the similarities between graph distributions. Table 6 presents the computed Cosine similarity and Euclidean distance between the distribution embeddings of the original graph h, the ground truth explanation subgraph $\textit{{h}}^{\prime}$, and the generated proxy graph $\tilde{\textit{{h}}}$. Notably, our proxy graph embedding $\tilde{\textit{{h}}}$ exhibits higher Cosine similarity scores and lower Euclidean distance with the original graph embedding h compared to the ground truth explanation embedding $\textit{{h}}^{\prime}$. We observe an average improvement of $17.9\%$ in Cosine similarity and $33.6\%$ in Euclidean distance. Particularly in the BA-2motifs dataset, there is a significant improvement of $60.4\%$ in Cosine similarity and $51.8\%$ in Euclidean distance. These findings underscore the effectiveness of our ProxyExplainer method in mitigating distribution shifts caused by induction bias in the to-be-explained GNN model $f(\cdot)$, thereby enhancing explanation performance.
Table 6: The Cosine score and Euclidean distance between the distribution embedding of the original graph h, explanation subgraph $\textit{{h}}^{\prime}$, and our proxy graph $\tilde{\textit{{h}}}$ on different datasets.
BA-2motifs
BA-3motifs
Alkane-Carbonyl
Benzene
Fluoride-Carbonyl
MUTAG
Avg. Cosine(h, $\textit{{h}}^{\prime}$) $\uparrow$
0.571
0.686
0.879
0.859
0.906
0.883
Avg. Cosine(h, $\tilde{\textit{{h}}}$) $\uparrow$
0.916 0.918 0.938 0.905 0.908 0.985
Avg. Euclidean(h, $\textit{{h}}^{\prime}$) $\downarrow$
1.210
1.199
0.912
0.893
0.805
0.975
Avg. Euclidean(h, $\tilde{\textit{{h}}}$) $\downarrow$
0.583 0.613 0.719 0.767 0.779 0.368
@@ D.2 Extensive Ablation Study
((d)) Benzene
((e)) Fluoride-Carbonyl
((f)) MUTAG
@@ D.3 Extensive experiment with GIN
In order to show the robustness of our ProxyExplainer in explaining different GNN models, we first pre-train GIN on two synthetic datasets and two real-world datasets to ensure that GIN has the ability to classify graphs accurately. The results are displayed in Figure 7. Then we use both baseline methods in the previous experiment and our method ProxyExplainer to provide explanations for the pre-trained GIN model. As seen in Figure 8, it is noticeable that ProxyExplainer achieves the best performance on these datasets. Specifically, it improves the AUC scores of $45.9\%$ on BA-2motifs, $6.1\%$ on BA-3motifs, $2.4\%$ on Alkane-Carbonyl, and $12.5\%$ on Fluoride-Carbonyl. From the analysis, we can see that our ProxyExplainer can identify accurate explanations across different datasets.
@@
((a)) BA-2motifs
((b)) BA-3motifs
((c)) Alkane-Carbonyl
((d)) Fluoride-Carbonyl
@@ D.4 Parameter Sensitive Analysis
Figure 9 shows the performance of our ProxyExplainer with respect to $\lambda$ on two synthetic datasets (BA-2motifs and BA-3motifs) and two real-world datasets (Alkane-Carbonyl and MUTAG). From Figure 9, we can see that ProxyExplainer consistently outperforms the best baseline, MixupExplainer, for $\lambda\in[0.25,1.0]$. This indicates that ProxyExplainer is stable. Figure 10 presents how the dimension of node latent embedding affects performance in ProxyExplainer, the results show that ProxyExplainer can reach the best performance at $D=512$.
((a)) BA-2motifs
((b)) BA-3motifs
((c)) Alkane-Carbonyl
((d)) MUTAG
@@
((a)) BA-2motifs
((b)) BA-3motifs
((c)) Alkane-Carbonyl
((d)) MUTAG

### v2

@@ 6.4 Ablation Studies (RQ3)
Figure 4: Visualization of explanation results from different explanation models on Benzene. The generated explanations are highlighted with bold orange edges.
@@
Figure 5: Visualization of explanation results from different explanation models on BA-2motifs. The generated explanations are highlighted with bold black edges.
@@ 6.5 Case Studies
In this part, we conduct case studies to qualitatively compare the performances of ProxyExplainer against others. We adopt examples from Benzene and BA-2motifs in this part. We show visualization results in Figure 4 and Figure 5. Explanations are highlighted with bold black and bold orange edges, respectively. From the results, our ProxyExplainer stands out by generating more compelling explanations compared to baselines. Specifically, ProxyExplainer maintains clarity without introducing irrelevant edges and exhibits more concise results compared to alternative methods. The visualized performance underscores ProxyExplainer’s ability to provide meaningful and focused subgraph explanations. More visualizations on these two datasets can be found in the Appendix D.5.
@@ 7 Conclusion
Acknowledgments
This project was partially supported by NSF grant IIS-2331908. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ Appendix A Notations
Symbols Descriptions
$\mathcal{G}$ A set of graphs
$G,{\mathcal{V}},{\mathcal{E}}$ Graph instance, node set, edge set
$v_{i}$ The $i$-th node
${\bm{X}}$ Node feature matrix
${\bm{A}}$ Adjacency matrix
${\bm{Z}}$ Node representation matrix
$Y$ Label of graph $G$
$\mathcal{Y}$ A set of labels
$G^{*}$ Optimal explanatory subgraph
${\mathcal{G}}^{*}$ A set of $G^{*}$
$G^{\prime}$ Candidate explanatory subgraph
$G^{\Delta}$ Non-explanatory graph
$\tilde{G}$ Proxy graph of $G^{\prime}$ with a fixed distribution
h, $\textit{{h}}^{\prime}$, $\tilde{\textit{{h}}}$ Graph embeddings
$d$ Dimension of node feature
$f(\cdot)$ To-be-explained GNN model
$\Psi_{\psi}(\cdot)$ Explanation function
$\psi$ Parameter of the explanation function
$P_{\mathcal{G}}$ Distribution of original training graphs
$P_{{\mathcal{G}}^{\prime}}$ Distribution of explanation subgraphs
$Q_{\bm{\phi}}$ Parameterized function of $P(\tilde{G}|G^{\prime})$
$\bm{\phi}$ Model parameters of $Q_{\bm{\phi}}$
$\bm{\phi}^{*}$ Optimal $\bm{\phi}$
$\alpha$ Balance parameter between $I(G,G^{\prime})$ and $I(Y,G^{\prime})$
$\bar{{\mathcal{E}}}$ The set of node pairs that are unconnected in $G$
$\tilde{p}_{uv}$ Probability of node pair $(u,v)$ in $\tilde{G}$
$\beta$ A hyper-parameter to get a trade-off between connected and unconnected node pairs
$f_{\text{enc}}(\cdot)$ The front part of GNNs that learns node representations
$f_{\text{cls}}(\cdot)$ The back part of GNNs that predicts graph labels based on node embeddings
$\sigma(\cdot)$ Sigmoid function
$\tau$ Temperature hyper-parameter for approximation
$\lambda$ A hyper-parameter in Proxy loss function
$\mathcal{L}_{\text{dist}}$ Distribution loss between $\tilde{G}$ and $G$
$\mathcal{L}_{\text{KL}}$ KL divergence between distribution of ${\bm{Z}}^{\Delta}$ and its prior
$\mathcal{L}_{\text{proxy}}$ Proxy loss
$\mathcal{L}_{\text{exp}}$ Explainer loss
@@ Appendix B Algorithm
           Update parameters in the explainer with backpropagation
@@ Appendix C Full Experimental Setups
We conduct all experiments on a Linux machine with 8 Nvidia A100-PCIE GPUs, each with 40GB of memory. The CUDA version is 12.4 and the driver version is 550.54.15. We use Python 3.9 and Torch 2.0.1 in our project. The code is available at https://github.com/realMoana/ProxyExplainer.
@@ C.1 Datasets
Benzene (Sanchez-Lengeling et al., 2020). Benzene consists 12,000 molecular graphs from the ZINC15(Sterling & Irwin, 2015) database, which can be classified into two classes. The main goal is to determine if a Benzene ring is exited in each molecule. In cases where multiple Benzene rings are present, each Benzene ring is treated as a distinct explanation.
Alkane-Carbonyl (Sanchez-Lengeling et al., 2020). The Alkane-Carbonyl dataset consists of a total of 4,326 molecule graphs that are categorized into two different classes. Positive samples refer to molecules that have both alkane and carbonyl functional groups. The ground-truth explanation includes alkane and carbonyl functional groups in a given molecule.
Fluoride-Carbonyl (Sanchez-Lengeling et al., 2020). The Fluoride-Carbonyl dataset has 8,671 molecular graphs. The ground-truth explanation is based on the particular combination of fluoride atoms and carbonyl functional groups present in each molecule.
@@
BA-3motifs (Chen et al., 2023b). BA-3motifs is an extended dataset inspired by the BA-2motifs and contains 3,000 synthetic graphs. Each base graph is accompanied by one of three different patterns: house, cycle, or grid.
Table 6: Statistics of molecule datasets for graph classification with ground-truth explanations.
MUTAG Benzene Alkane-Carbonyl Fluoride-Carbonyl BA-2motifs BA-3motifs
Graphs 4,337 12,000 4,326 8,671 1,000 3,000
Average nodes 29.15 20.58 21.13 21.36 25.00 21.92
Average edges 60.83 43.65 44.95 45.37 25.48 29.51
Node features 14 14 14 14 10 4
Original graph [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground truth explanation [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
NH2, NO2 Benzene Ring Alkane,C=O F-,C=O House, cycle House,cycle, grid
@@
MUTAG Benzene Alkane-Carbonyl Fluoride-Carbonyl BA-2motifs BA-3motifs
Train Acc 0.850 0.930 0.979 0.951 0.999 0.997
Val Acc 0.834 0.927 0.986 0.956 1.0 0.997
Test Acc 0.804 0.915 0.975 0.951 1.0 0.977
@@ C.2 Baselines
To evaluate our model, we well-train the GCN model in to ensure that it takes good performance in graph classification tasks. The results are displayed in Table 7. For a comprehensive comparison, we incorporate various post-hoc explanation methods, including GradCAM, GNNExplainer, PGExplainer, ReFine, and MixupExplainer.
@@ D.1 Extensive Distribution Analysis
As shown in previous work (Zhang et al., 2023b), another way to approximate the distribution differences of graphs is to compare their vector embeddings. Here, we adopt the same setting and use Cosine similarity and Euclidean distance to intuitively approximate the similarities between graph distributions. Table 8 presents the computed Cosine similarity and Euclidean distance between the distribution embeddings of the original graph h, the ground truth explanation subgraph $\textit{{h}}^{\prime}$, and the generated proxy graph $\tilde{\textit{{h}}}$. Notably, our proxy graph embedding $\tilde{\textit{{h}}}$ exhibits higher Cosine similarity scores and lower Euclidean distance with the original graph embedding h compared to the ground truth explanation embedding $\textit{{h}}^{\prime}$. We observe an average improvement of $19.5\%$ in Cosine similarity and $35.6\%$ in Euclidean distance. Particularly in the BA-2motifs dataset, there is a significant improvement of $60.4\%$ in Cosine similarity and $51.8\%$ in Euclidean distance. These findings underscore the effectiveness of our ProxyExplainer method in mitigating distribution shifts caused by induction bias in the to-be-explained GNN model $f(\cdot)$, thereby enhancing explanation performance.
Table 8: The Cosine similarity score and Euclidean distance between the distribution embeddings of the original graph h, explanation subgraph $\textit{{h}}^{\prime}$, and our proxy graph $\tilde{\textit{{h}}}$ on different datasets.
MUTAG Benzene Alkane-Carbonyl Fluoride-Carbonyl BA-2motifs BA-3motifs
Avg. Cosine(h, $\textit{{h}}^{\prime}$) $\uparrow$ 0.883 0.835 0.889 0.904 0.571 0.686
Avg. Cosine(h, $\tilde{\textit{{h}}}$) $\uparrow$ 0.985 0.905 0.938 0.908 0.916 0.918
Avg. Euclidean(h, $\textit{{h}}^{\prime}$) $\downarrow$ 0.975 1.010 0.940 0.806 1.210 1.199
Avg. Euclidean(h, $\tilde{\textit{{h}}}$) $\downarrow$ 0.368 0.767 0.719 0.779 0.583 0.613
@@ D.2 Extensive Ablation Study
((d)) Fluoride-Carbonyl
((e)) BA-2motifs
((f)) BA-3motifs
@@ D.3 Extensive Experiment with GIN
In order to show the robustness of our ProxyExplainer in explaining different GNN models, we first pre-train GIN on two real-world datasets and two synthetic datasets to ensure that GIN has the ability to classify graphs accurately. The results are displayed in Figure 7. Then we use both baseline methods in the previous experiment and our method ProxyExplainer to provide explanations for the pre-trained GIN model. As seen in Figure 8, it is noticeable that ProxyExplainer achieves the best performance on these datasets. Specifically, it improves the AUC scores of $2.9\%$ on Alkane-Carbonyl, $7.9\%$ on Fluoride-Carbonyl, $45.9\%$ on BA-2motifs, and $6.1\%$ on BA-3motifs. From the analysis, we can see that our ProxyExplainer can identify accurate explanations across different datasets.
@@
((a)) Alkane-Carbonyl
((b)) Fluoride-Carbonyl
((c)) BA-2motifs
((d)) BA-3motifs
@@ D.4 Parameter Sensitive Analysis
Figure 9 shows the performance of our ProxyExplainer with respect to $\lambda$ on two real-world datasets (MUTAG and Alkane-Carbonyl) and two synthetic datasets (BA-2motifs and BA-3motifs). From Figure 9, we can see that ProxyExplainer consistently outperforms the best baseline, MixupExplainer, for $\lambda\in[0.25,1.0]$. This indicates that ProxyExplainer is stable. Figure 10 presents how the dimension of node latent embedding affects performance in ProxyExplainer, the results show that ProxyExplainer can reach the best performance at $D=512$.
((a)) MUTAG
((b)) Alkane-Carbonyl
((c)) BA-2motifs
((d)) BA-3motifs
@@
((a)) MUTAG
((b)) Alkane-Carbonyl
((c)) BA-2motifs
((d)) BA-3motifs